# Elastic, Quasielastic, and Superelastic Electron Scattering from Thermal Lattice Distortions in Perfect Crystals

### v1

@@ Abstract
In conventional treatments of electron transport, momentum relaxation within a perfect, defect-free crystal is commonly assumed to require phonon creation or annihilation. Here we treat the crystal as finite and isolated, retaining the lattice center-of-mass (recoil) degree of freedom and enforcing conservation of total mechanical momentum alongside discrete crystal pseudomomentum. Starting from the density-density form of the electron-lattice interaction, we find that an electron in the interior of a perfect crystal admits strong, and in some regimes dominant, elastic momentum-relaxing scattering channels, in which momentum is conserved by recoil of the lattice background without phonon excitation. In addition, we identify mixed quasielastic and superelastic channels in which phonon occupations change but do not account fully for the electron’s momentum transfer. These results provide a microscopic basis for momentum relaxation that does not rely on local energy dissipation. They naturally reconcile a wide range of experimental observations—including weak localization, quantum oscillations, ultrasonic attenuation, and the separation of momentum and energy relaxation times—with predominantly elastic scattering in clean crystals. The framework clarifies how diffusive transport, including Planckian-scale diffusion, can emerge from elastic dynamics in a time-dependent lattice background.
@@ I Introduction
Electron-lattice interactions are among the most important processes in condensed matter physics. They govern momentum relaxation, energy dissipation, and the loss of quantum coherence that underlies electrical resistance. In conventional transport theory, electron momentum relaxation in a perfect, defect-free crystal is commonly identified with inelastic scattering processes in which phonons are created or annihilated. Within this view, deflection of an electron by thermal lattice motion is inseparable from energy transfer to internal vibrational degrees of freedom.
This perspective is codified in the Bloch-Grüneisen and Boltzmann approaches to resistivity, which treat electronic transport as an incoherent sequence of scattering events and assume no electronic phase coherence beyond that retained in lowest-order scattering matrix elements. It is further reinforced by the Fröhlich interaction [10], which couples electrons explicitly to phonon creation and annihilation operators and therefore enforces inelasticity at first order in perturbation theory. Within this canonical framework, thermal deflection of electrons in a perfect crystal is taken to require phonon excitation, while elastic scattering is associated primarily with defects or impurities [39, 4].
Here we present evidence that this identification is overly restrictive. We show that thermally induced electron scattering in a pure crystal is often largely elastic when viewed on an event-by-event basis. Elastic and quasi-elastic momentum-relaxing channels exist even in a perfectly periodic lattice, without the necessity of phonon creation or annihilation. Previously, elastic scattering was almost exclusively associated with explicit disorder; we demonstrate that this association is not required by symmetry, conservation laws, or kinematics.
Early discussions of electronic transport already contained hints of this distinction. Peierls emphasized that electrical resistance requires degradation of electronic (crystal) momentum rather than dissipation of electronic energy, and noted explicitly that the lattice, as a massive collective object, can absorb momentum without requiring significant energy transfer [27]. Pippard similarly argued that defects and impurities are important but not indispensable, and that momentum relaxation could, in principle, proceed through elastic or quasi-elastic mechanisms even in very pure crystals [29]. These insights acknowledged the logical separation between momentum relaxation and energy dissipation, although this separation was not consistently maintained in subsequent formulations of transport theory.
In contrast, the canonical Fröhlich Hamiltonian and Bloch-Grüneisen framework [7] implicitly exclude rigid translation of the crystal as a whole from the dynamical description. The lattice center-of-mass (recoil) degree of freedom is effectively clamped, and momentum transfer is assigned exclusively to internal phonon excitations or to Umklapp processes. Once these assumptions are imposed, momentum loss is forced to coincide with energy dissipation, and elastic deflection by a perfect lattice is excluded by construction.
We revisit this assumption by treating the crystal as finite, isolated, and freely floating, and by explicitly retaining its total mechanical momentum. Enforcing conservation of total momentum more completely cannot be formally incorrect; at worst it may prove quantitatively unimportant. As we show below, however, restoring the lattice center-of-mass degree of freedom qualitatively alters the available scattering channels for an electron deep inside a pure conducting crystal. In particular, it reveals an elastic recoil channel that is absent by design in the conventional formulation.
The physical content of this channel is elementary. If a particle embedded in a composite object receives a sudden impulse, the momentum of the whole must change immediately, even if the internal degrees of freedom remain largely unexcited. A familiar example is provided by Mössbauer scattering, where momentum is absorbed predominantly by recoil of the entire lattice while phonon excitation is strongly suppressed. As we shall show, an analogous mechanism operates for electrons interacting with a thermally fluctuating but otherwise perfect crystal.
The ambiguity of elastic scattering
Elastic scattering is most familiar in the context of impurity or defect scattering. In that setting, momentum conservation is usually justified in one of two ways. First, the presence of an impurity breaks discrete translational symmetry, so crystal pseudomomentum need not be conserved. Second, within a particle picture, one argues that if an electron enters with one momentum and exits with another, the atomic lattice must recoil so as to conserve total mechanical momentum. Both statements are essentially correct. However, neither addresses the internal bookkeeping problem that arises for an electron scattering deep inside a perfectly periodic crystal: how does a change in the electron’s pseudomomentum communicate with the mechanical momentum of the lattice as a whole?
This unresolved point lies at the heart of the transport problem in a translationally invariant system. As we show below, it is resolved once three elements are made explicit: (i) the lattice is treated as a fully dynamical object rather than a fixed reference, (ii) pseudomomentum is assigned not only to the electron but also to the crystal background in the absence of the electron density (see the next section), and (iii) the electron-lattice interaction is formulated in density-density form [24, 10], which makes the exchange of pseudomomentum between electron and lattice explicit even in a perfectly periodic crystal.
In any untethered collection of atoms, an impulse imparted to even a single constituent instantaneously changes the total momentum of the assembly. While the subsequent redistribution of that momentum among internal degrees of freedom proceeds causally and at finite velocity, the change in the collective momentum itself is immediate. This total momentum is not an abstract bookkeeping device but a genuine dynamical variable that appears explicitly in the many-body wavefunction.
In the present formulation, this momentum conservation is enforced transparently. As shown in Eq. 12, a momentum transfer $\mathbf{q}$ to the foreground electron is accompanied by an equal and opposite momentum $\mathbf{q}$ carried by the remaining degrees of freedom, which we identify as the lattice background. This statement holds independently of whether phonons are created or destroyed in the process.
@@
Finally, we clarify our use of the term elastic. In a finite-temperature many-body system, strictly elastic processes do not exist in an absolute sense. Even the dominant “elastic” line in Mössbauer scattering exhibits slight thermal broadening, reflecting the thermal motion of the emitting nucleus and its finite velocity autocorrelation time. The term elastic is nevertheless standard in this context, and we adopt it here with the same understood margin. When we use the term quasielastic, we instead refer to processes involving energy exchange that is appreciably larger than the thermal broadening scale, but still too small to account for the full momentum change of the electron.
@@ II Foreground, Background, and Pseudomomentum
The conceptual distinction between momentum, crystal momentum, and electrical current has been recognized since the early development of band theory, but its implications for microscopic transport have remained subtle. In a particularly prescient analysis, Kohn emphasized that crystal momentum is a conserved quantity associated with translational symmetry, while the electrical current need not be proportional to it, and that momentum may be exchanged with the lattice without the necessity of phonon excitation or energy dissipation [17]. This observation already points to the possibility that momentum conservation can be satisfied through coherent motion of the lattice background rather than through inelastic scattering processes. What has remained largely implicit, however, is how this distinction should be implemented at the level of a local scattering event occurring deep in the interior of a crystal. In particular, the possibility that total crystal may be explicitly partitioned between a tagged electron and the remaining lattice, both with pseudomomenta appearing as dynamical variables in the many-body wavefunction, has not been systematically developed.
The present work makes this partition explicit and operational, thereby extending Kohn’s insight to a microscopic framework in which elastic deflection, coherent lattice recoil, and diffusive transport can coexist naturally in a perfect crystal. In what follows, we formalize a symmetric foreground-background structure, derive the elastic scattering amplitude within first-order perturbation theory, and compute the elastic fraction explicitly. We show that this fraction is close to unity for copper at room temperature. We also clarify how the present formulation extends the Fröhlich framework without contradicting its successful predictions, and discuss the implications for the interpretation of transport phenomena in crystalline solids.
@@
In a crystalline environment, the electronic Hamiltonian is invariant under discrete translations of the electron by lattice vectors, with the background held fixed. The associated conserved quantity is crystal momentum (pseudomomentum); for an electron in a periodic potential this is the familiar Bloch momentum $\hbar\mathbf{k}$.
@@
Figure 1: Sliding Blocks. Floating in space, two blocks attract one another and slide frictionlessly along their common long axis, $x$. At the top, both blocks carry conserved mechanical momenta, and clearly $P_{\rm tot}=p_{A}+p_{B}$. At the bottom, the same relation holds, but $p_{A}$ and $p_{B}$ are now pseudomomenta; their sum nevertheless equals the conserved mechanical momentum. The two blocks serve as surrogates for a foreground electron (A) in a perfect background crystal (B), illustrating that both the foreground and the background acquire pseudomomentum in each other’s presence.
@@
Earlier we remarked that a pseudomomentum cannot be combined with a mechanical momentum. We have now seen, however, that two pseudomomenta can combine to form a mechanical momentum. These rules are analogous to those governing the collision of two objects in free space, except that here the interaction is periodic in their relative coordinate, apart from defects and lattice vibrations.
The total momentum $\hat{\mathbf{Q}}$ is conserved, but the foreground and background momenta, $\hat{\mathbf{p}}$ and $\hat{\mathbf{P}}$, undergo equal and opposite changes, in exact analogy with collisions in free space. This exchange does not require the creation or annihilation of internal lattice excitations and occurs immediately at the collision event (see Fig. 2).
When a crystal containing an internal electron is treated as a closed many-body system, the total background momentum $\mathbf{P}=\sum_{i}\mathbf{p}_{i}$ is defined at each instant as the generator of translations of the background degrees of freedom. Since the foreground is held fixed in this construction, $\mathbf{P}$ is a pseudomomentum. An impulse imparted to a single atom in the background by a foreground object changes the total background momentum immediately by the same amount, resulting in a shift of the center-of-mass velocity $\Delta\mathbf{V}_{\mathrm{CM}}=\Delta\mathbf{p}/M$, independent of any subsequent internal redistribution of momentum.
The finite-speed propagation of stress, phonons, or other internal excitations governs how momentum is shared among individual degrees of freedom, but it does not affect the instantaneous value of the conserved quantity itself. This kinematic distinction follows directly from translational invariance and does not imply nonlocal dynamics or superluminal signaling; rather, it reflects the separation between global conservation laws and the causal evolution of local observables.
@@
The interaction is periodic only over a finite distance $L$, as required for any physically realizable system. This situation is directly analogous to the classic particle in a box. Momentum is not strictly conserved in such a system; nevertheless, traveling wave packets carrying well-defined momentum can be constructed as linear combinations of standing-wave eigenstates, which individually carry no momentum. Over finite times these wave packets propagate as if momentum were conserved, until a collision with a boundary reverses their motion.
This type of approximate symmetry, valid only over finite spatial and temporal windows, is sometimes implicit in discussions of wave-packet dynamics in finite systems. For clarity, we refer to it here as finite-window symmetry. If one insists on working with finite crystals, this concept must be adopted: translational symmetry holds only within a spatial and temporal window set by the system size.
In a large crystal this construction is an excellent approximation. Localized corpuscles can be formed that propagate nearly as Bloch waves with well-defined pseudomomentum and experience long travel times before encountering a boundary. We emphasize that this corpuscle construction is a calculational convenience that permits the use of the notation and results associated with the formally unphysical Born-von Kármán boundary conditions, without assuming their literal validity.
@@ III Lattice degrees of freedom and zero modes
The classical and quantum theory of lattice vibrations explicitly separates rigid motion from internal vibrations. Born and Huang showed that the equations of motion of a finite crystal admit three zero-frequency normal modes corresponding to rigid translation of the crystal as a whole, while the internal normal modes satisfy the mass-weighted constraint $\sum_{j}m_{j}\hat{\mathbf{u}}_{j}=\mathbf{0},$ so that internal vibrations carry no net mechanical momentum [8]. All mechanical momentum of the crystal resides in the translational zero mode.
The total center-of-mass degree of freedom of the combined foreground-background system is always conserved and need not be carried explicitly. For an interior electron, it is crucial to track the pseudomomentum of the foreground electron and separately the background,
@@ IV.1 Foreground and background Hamiltonians
Given our aim of formulating transport without assuming pre-existing quasiparticles, care is required in defining the Hamiltonian. We write the total Hamiltonian as
@@
This identity expresses the fact that a uniform translation of the lattice multiplies all density Fourier components by the phase $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, while all information about internal lattice dynamics is contained in $\hat{\rho}^{\rm int}_{\mathbf{q}}$.
@@
With this factorization, the electron–lattice interaction may be written in density–density form as
@@
which follows the standard Born-approximation scattering formalism introduced by Van Hove [34]. Using Eq. (10), this becomes
$H_{\rm int}$ $=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i\mathbf{q}\cdot\hat{\mathbf{r}}}\,e^{i\mathbf{q}\cdot\hat{\mathbf{R}}}\,\hat{\rho}^{\rm int}_{-\mathbf{q}}$ (12)
@@
Crucially, this exchange of pseudomomentum applies whether or not there is a change a change in phonon occupation numbers. Elastic matrix elements of $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ are generically nonzero. The background zero mode is therefore capable of absorbing finite pseudomomentum with negligible energy cost, providing the kinematic basis for elastic but momentum-relaxing scattering. Any combination of phonon involment and lattice background recoil is allowed. It seems this should matter to measureable properties.

### v2

@@ Abstract
In standard treatments of electron transport, momentum relaxation in a perfect, defect-free crystal is linked with phonon creation or annihilation. In this work, we reconsider this problem for a finite, isolated crystal, retaining the lattice center-of-mass (recoil) degree of freedom and enforcing conservation of total mechanical momentum together with discrete crystal pseudomomentum. Starting from the density–density form of the electron–lattice interaction, we show that an electron in the interior of a perfect crystal admits elastic momentum-transfer channels in which total momentum is conserved by recoil of the lattice background without phonon excitation. These elastic channels can provide the leading contribution to momentum relaxation. We further identify mixed quasielastic and superelastic processes in which phonon occupations change but do not account entirely for the electron’s momentum transfer. The elastic channels arise within the standard microscopic Hamiltonian and do not require additional disorder or defects. The resulting framework provides a complementary microscopic perspective on momentum relaxation in clean crystals, and is consistent with experimental phenomena such as weak localization, quantum oscillations, ultrasonic attenuation, and the observed separation of momentum and energy relaxation times.
@@ I Introduction
Electron–lattice interactions play a central role in condensed matter physics. They govern both momentum relaxation and energy exchange in crystalline solids, and therefore underlie electrical resistivity and the loss of electronic coherence. Experimentally, however, momentum and energy relaxation often occur on markedly different timescales. In clean metals over broad temperature ranges, electrical resistivity and magnetotransport measurements indicate rapid momentum randomization, while hot-electron experiments, Johnson-noise thermometry, and related nonequilibrium probes show that energy relaxation to the lattice can remain comparatively slow. The resulting hierarchy
$\tau_{p}\ll\tau_{E},$ (1)
where $\tau_{p}$ is the momentum-relaxation time and $\tau_{E}$ the energy-relaxation time, indicates that electronic momentum can be efficiently scrambled without substantial net energy transfer to internal lattice excitations.
In conventional transport theory, momentum relaxation in a perfect, defect-free crystal is typically described in terms of phonon-mediated scattering processes. Within the Bloch–Grüneisen and semiclassical Boltzmann frameworks, transport is treated as a sequence of transitions between electronic states, with rates evaluated using the electron–phonon interaction [35, 3]. In the standard Fröhlich Hamiltonian [9], the interaction is expressed explicitly in terms of phonon creation and annihilation operators, and momentum transfer is therefore associated, at lowest order, with changes in phonon occupation. In this commonly employed description, elastic momentum relaxation is primarily attributed to impurities or other forms of explicit disorder.
The observed separation between $\tau_{p}$ and $\tau_{E}$ motivates a careful examination of how momentum conservation is implemented microscopically in a perfectly periodic crystal. The conceptual distinction between momentum relaxation and energy dissipation has long been recognized. Peierls emphasized that electrical resistance requires degradation of electronic (crystal) momentum rather than dissipation of electronic energy, and noted that the lattice, as a massive collective object, can absorb momentum with minimal energy cost [24]. Pippard similarly stressed that while defects often play an important role, momentum relaxation need not be uniquely tied to strongly inelastic processes [25].
In this work we revisit the problem starting from the same microscopic electron–lattice Hamiltonian, but retaining explicitly the lattice center-of-mass degree of freedom. In conventional treatments the crystal is effectively considered as a fixed reference frame, and the translational zero mode does not appear among the dynamical variables. Here we adopt an expanded viewpoint in which the crystal total mechanical momentum is retained. This does not modify the underlying Hamiltonian; rather, it makes explicit a degree of freedom that is commonly suppressed.
When total momentum conservation is enforced in this manner, a transfer of momentum $\mathbf{q}$ to an electron is accompanied by an equal and opposite momentum $\mathbf{q}$ carried by the remaining degrees of freedom, which we identify as the lattice background. This exchange can occur without excitation of internal phonon modes, leading to an elastic recoil channel even in a perfectly periodic crystal. We also identify mixed quasielastic and superelastic processes in which phonon occupations change but do not fully account for the electron’s momentum transfer. These channels arise within the standard density–density form of the electron–lattice interaction [22, 9] and require no additional disorder.
The physical content of the elastic recoil channel is elementary. In any untethered composite object, an impulse imparted to one constituent immediately changes the total momentum of the whole. While redistribution of that momentum among internal degrees of freedom proceeds causally and at finite velocity, the collective momentum of the backgroud crystal appears in the wavefunction and is not dependent of the details of redistribution of the momentum among the constitutuents of the whole. . An instructive analogue is provided by Mössbauer scattering, in which the dominant elastic line is recoil of the entire lattice without internal phonon excitation.
A conceptual roadblock has been that, in contrast to processes such as Mössbauer and neutron scattering, where elastic, non-Bragg diffuse deflection is widely acknowledged, conduction electrons do not enter and exit; they start and stay in the crystal. The roadblock is addressed by assigning pseudomomentum to both a “foreground” electron density and a “background” consisting of everything else. We hasten to point out that the question here involves real physics, and not mere representation: Must phonons be created or destroyed when electrons deflect in a pure crystal, or not?
On the meaning of elastic scattering
Elastic scattering of interior electrons is most familiar in the context of impurity or defect scattering, where translational symmetry is explicitly broken. Here, we restore pseudomomentum to the status of a conserved momentum, and as in free space what then is conserved in such an elastic deflection? The answer will not be controversial: the total momentum of the whole system is conserved. What we believe is new is that two pseudmomomenta have changed equally and oppositely in such a collision: one belonging to the electron, another to the rest of the crystal. The total momentum, the sum of the two, is unchanged. In a perfectly periodic crystal, the relationship between changes in electronic pseudomomentum and the mechanical momentum of the lattice background is less often made explicit. In the present formulation this bookkeeping becomes transparent once (i) the lattice is treated as fully dynamical, (ii) pseudomomentum is assigned both to the electron and to the background in the absence of the electron density, and (iii) the interaction is expressed in density–density form, making the exchange of pseudomomentum explicit.
@@
Finally, we clarify our terminology. In a finite-temperature many-body system, strictly elastic processes do not exist in an absolute sense. Even the dominant line in Mössbauer scattering exhibits slight thermal broadening. We therefore use the term elastic to denote processes in which phonon occupation does not change, and quasielastic to denote processes involving energy exchange that is small compared with the momentum transfer involved.
@@ II Foreground, Background, and Pseudomomentum
The conceptual distinction between momentum, crystal momentum, and electrical current has been recognized since the early development of band theory, but its implications for microscopic transport have remained subtle. In a particularly prescient analysis, Kohn emphasized that crystal momentum is a conserved quantity associated with translational symmetry, while the electrical current need not be proportional to it, and that momentum may be exchanged with the lattice without the necessity of phonon excitation or energy dissipation [17]. This observation already points to the possibility that momentum conservation can be satisfied through recoil of the lattice background rather than through inelastic scattering processes. What has remained largely implicit, however, is how this distinction should be implemented at the level of a local scattering event occurring deep in the interior of a crystal. Here, we explicitly assign pseudomomentum to both a foreground electron and the background lattice, with both pseudomomenta appearing as dynamical variables in the many-body wavefunction.
Making this partition explicit is a realization of Kohn’s insight, within a framework in which elastic deflection, coherent lattice recoil, and diffusive transport can coexist naturally in a perfect crystal. In what follows, we formalize a symmetric foreground-background structure, derive the elastic scattering amplitude within first-order perturbation theory, and compute the elastic fraction explicitly. We show that this fraction is close to unity for copper at room temperature. We also clarify how the present formulation extends the Fröhlich framework without contradicting its successful predictions, and discuss the implications for the interpretation of transport phenomena in crystalline solids.
@@
In a crystalline environment, the electronic Hamiltonian is invariant under discrete translations of the electron by lattice vectors, with the background held fixed. The associated conserved quantity is pseudomomentum (crystal momentum); for an electron in a periodic potential this is the familiar Bloch momentum $\hbar\mathbf{k}$.
@@
Figure 1: Sliding Blocks. Floating in space, two blocks attract one another and slide frictionlessly along their common long axis. At the top, both blocks carry conserved mechanical momenta, and clearly $P_{\rm tot}=p_{A}+p_{B}$. At the bottom, the same relation holds, but $p_{A}$ and $p_{B}$ are now pseudomomenta; their sum nevertheless equals the conserved mechanical momentum. The two blocks serve as surrogates for a foreground electron (A) in a perfect background crystal (B), illustrating that both the foreground and the background acquire pseudomomentum in each other’s presence.
@@
Even under the usual rule stating that defects, such as a crystal edge or impurity, scatter electrons elastically, there is a subtlety: the electron changes its pseudomomentum, but the lattice cannot the respond or recoil in kind. A pseudomomentum cannot be combined with a mechanical momentum. We have now seen, however, that both foreground and background have pseudomomentum, and can exchage it freely. Total momentum is conserved for example in the very real deflection of an electron at an edge by equal and opposite pseudomomentum changes.
These rules are analogous to those governing the collision of two objects in free space, except that here the interaction is periodic in their relative coordinate, apart from defects and lattice vibrations.
The total momentum $\hat{\mathbf{Q}}$ is conserved, but the foreground and background momenta, $\hat{\mathbf{p}}$ and $\hat{\mathbf{P}}$, undergo equal and opposite changes, in exact analogy with collisions in free space. This exchange does not require the creation or annihilation of internal lattice excitations and occurs immediately at the collision event (see Fig. 2). The very real recoil of the atom labeled in red in Fig. 2 may very well prove to have excited no vibrational quanta when measured in a number state basis. This is just the elementary fact that a slightly displaced ground state of a harmonic oscillator is still mostly in the ground state, if measured.
Two measurable subsystem momenta and the total momentum
Figure 1 shows two variations on a mechanical system. At the top, two smooth blocks interact frictionlessly. At the bottom, both blocks are corrugated, leading to periodic interaction.
@@
In any physically realizable crystal the interaction is periodic only over a finite spatial extent $L$. Strict translational invariance is therefore absent at the boundaries, and crystal momentum is not an exact conserved quantity for the full finite system. This situation is directly analogous to the textbook particle in a box: the exact eigenstates are standing waves that do not carry definite momentum, yet traveling wave packets with well-defined average momentum can be constructed as superpositions of these eigenstates. Over times short compared with the traversal time to the boundary, such wave packets propagate as if momentum were conserved.
The same reasoning applies to a large but finite crystal. Although global translational symmetry is broken by the boundaries, the Hamiltonian is locally periodic in the bulk. Wave packets localized far from the boundaries evolve, for long but finite times, according to the translationally invariant bulk Hamiltonian. In this regime one may assign a well-defined crystal pseudomomentum to the packet, up to corrections that vanish as $L\to\infty$.
This type of approximate symmetry is standard in discussions of the thermodynamic limit, bulk transport theory, and wave-packet dynamics in finite systems. For clarity, we refer to it here as finite-window symmetry: translational invariance holds within a spatial and temporal window set by the system size. No new symmetry is implied; rather, this terminology emphasizes that the use of bulk momentum quantum numbers in a finite crystal is an approximation controlled by the ratio of the observation time and length scales to the system size.
In a macroscopic crystal this approximation is exceedingly accurate. Localized wave packets (“corpuscles”) can be constructed that propagate nearly as Bloch waves with well-defined pseudomomentum and experience long travel times before interacting with boundaries. This construction permits the use of the notation and results associated with Born–von Kármán boundary conditions as a calculational device, without assuming literal periodicity of the finite sample.
In a finite crystal, the exact normal modes are standing waves. As exact eigenmodes they have zero net mechanical momentum. Any “phonon carrying momentum $\hbar q$” is then not an eigenstate; it’s a wavepacket (a superposition of standing modes) that only behaves like a traveling object transiently until it reflects, mode-mixes, etc.
@@ III Lattice degrees of freedom and zero modes
The classical and quantum theory of lattice vibrations explicitly separates rigid translation from internal vibrations. Born and Huang showed that the equations of motion of a finite crystal admit three zero-frequency normal modes corresponding to rigid translation of the crystal as a whole, while the internal normal modes satisfy the mass-weighted constraint $\sum_{j}m_{j}\hat{\mathbf{u}}_{j}=\mathbf{0},$ so that internal vibrations carry no net mechanical momentum [7]. All mechanical momentum of the crystal resides in the translational zero mode.
@@ IV.1 Foreground and background Hamiltonians
. We write the total Hamiltonian as
@@
Here $\hat{\mathbf{R}}$ is the lattice center-of-mass operator, $\bm{\xi}^{(0)}_{j}$ are equilibrium positions, and $\hat{\mathbf{u}}_{j}$ are internal displacements. Equation (11) is an exact operator identity, independent of any choice of basis (plane waves, Bloch waves, or standing waves).
This identity expresses a purely kinematic fact: a rigid translation of the lattice multiplies every density Fourier component by the phase $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$, while all internal lattice dynamics are contained entirely in $\hat{\rho}^{\rm int}_{\mathbf{q}}$.
@@
The electron–lattice interaction may be written in density–density form as
@@
following the formalism of Van Hove [30]. No assumption about plane-wave eigenstates is required; $\mathbf{q}$ labels Fourier components of the interaction kernel.
Using Eq. (11), this becomes
$H_{\rm int}$ $=\sum_{\mathbf{q}}v(\mathbf{q})\,e^{i{\mathbf{q}}\cdot\hat{\bm{\rho}}}\,\,\hat{\rho}^{\rm int}_{-\mathbf{q}},$ (13)
@@
Crucially, this exchange of pseudomomentum applies whether or not there is a change a change in phonon occupation numbers.
Elastic matrix elements of $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ are generically nonzero. The background zero mode is therefore capable of absorbing finite pseudomomentum with negligible energy cost, providing the kinematic basis for elastic but momentum-relaxing scattering. Any combination of phonon involvement and lattice background recoil is allowed. It seems this should matter for measurable properties.